mhchen:

Is this a valid step in math:

mhchen:

\[|a|\frac{\epsilon}{2(|a|+1)} < \frac{\epsilon}{2}\]

mhchen:

\[|a|\frac{\epsilon}{2|a|+2} < \frac{\epsilon}{2}\]

mhchen:

\[|a|\frac{\epsilon}{2|a|+2} < |a|\frac{\epsilon}{2|a|}\] right?

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mhchen \[|a|\frac{\epsilon}{2|a|+20000000} < |a|\frac{\epsilon}{2|a|}\] right? \(\color{#0cbb34}{\text{End of Quote}}\)

mhchen:

Huh...it is right.